# Weixiang Zhao1, Yulin Hu1   Jiahe Guo1footnotemark:    Xingyu Sui1footnotemark:   , Tongtong Wu2, Yang Deng3 Yanyan Zhao1, Bing Qin1, Wanxiang Che1, Ting Liu1 1

### v1

@@ 4.1 Experimental Setup
We select one representative English-centric LLMs: LLaMA-3-8B-Instruct (AI@Meta, 2024) and 2 MLLMs: LLaMA-3.1-8B-Instruct [^1] https://huggingface.co/meta-llama/Llama-3.1-8B-Instruct and Phi-3.5-mini-instruct (Abdin et al., 2024), to fully validate the effectiveness and generality of our Lens in enhancing multilingual performance.
@@
In the bilingual setting, English (En) serves as the central language, while Chinese (Zh) is chosen as the target language for expansion. Chinese is selected due to its growing prominence in the academic focus on multilingual enhancement for LLMs.
In the multilingual setting, we select six target languages for enhancement based on the availability of language resources. The high-resource languages are Chinese (Zh) and Japanese (Jp); the medium-resource languages are Korean (Ko) and Arabic (Ar); and the low-resource languages are Bengali (Bn) and Swahili (Sw), with English (En) continuing to serve as the central language.
It is important to note that these target languages are classified as out-of-scope in the official model card of the above LLMs and MLLMs, which further underscores their relevance for enhancement.
@@
We sample 300 data points from the Aya Dataset for each language to probe the language space and 200 data points from the Bactrian-X dataset per language to manipulate the language space.
@@
  • Multilingual Understanding: (1) XCOPA (Ponti et al., 2020), (2) XWinograd (Muennighoff et al., 2023), (3) XStoryCloze (Lin et al., 2022) and (4) M-MMLU (Hendrycks et al., 2021; Lai et al., 2023). Accuracy is adopted as the evaluation metric and we randomly sample up to 1,000 data points from each benchmark for evaluation.
  • Multilingual Generation: (5) MT-Bench (Zheng et al., 2023): The dataset is designed for open-ended generation to evaluate a model’s ability to follow multi-turn instructions. The evaluation follows the LLM-as-a-judge approach, where GPT-4o is prompted to assign a score directly to a single response on a scale of 1 to 10. It is essential to highlight that the languages targeted for enhancement, as mentioned above, are all within the capability range of GPT-4o, especially given that its official model card [^2] https://cdn.openai.com/gpt-4o-system-card.pdf. emphasizes support for low-resource languages such as Swahili (sw) and Bengali (bn). This underscores the validity and reliability of the evaluation approach. In addition, we employ Language Fidelity (Holtermann et al., 2024) as a metric to assess the consistency between input and output languages, offering a clear measure of how effectively different methods mitigate the model’s off-target issues.
@@ 4.2 Baseline Methods
For comparison, we consider the following baseline methods that enhance LLMs’ multilingual capabilities using multilingual instruction fine-tuning technique: (1) xSFT & xSFT-Full (Ouyang et al., 2022): xSFT performs multilingual instruction fine-tuning using the same data volume as our Lens. In contrast, xSFT-Full utilizes the full dataset for each target language from the Aya Collection and Bactrian-X. (2) QAlign (Zhu et al., 2024a): It explores the benefits of question alignment, where the model is trained to translate inputs into English by finetuning on X-English parallel question data. (3) SDRRL (Zhang et al., 2024c): It is based on self-distillation from resource-rich languages that effectively improve multilingual performance by leveraging self-distillated data.
Table 1: Detailed results on the multilingual understanding and multilingual generation benchmarks with the English-centric LLaMA-3-8B-Instruct backbone and the multilingual LLaMA-3.1-8B-Instruct backbone under the bilingual setting (English and Chinese). Accuracy serves as the evaluation metric for multilingual understanding, while GPT-4o ratings (on a scale of 1 to 10) are provided for MT-Bench. The values in parentheses represent language fidelity. Results highlighted in green indicate an improvement or performance comparable to the original backbone, while those highlighted in red signal a decline in performance relative to the original backbone.
@@ 4.3 Implementation Details
Lens is a model-agnostic multilingual enhancement method that is compatible with different transformer-based LLM. Our experiments are implemented with PyTorch (Paszke et al., 2019) and Transformer library (Wolf et al., 2020) on a single NVIDIA A800-SXM4-80GB GPU. The training duration is set to one epoch with the learning rate of 1e-5 and batch size of 8 across all backbones. For more detailed settings, please refer to the Appendix D.
@@ 4.4 Overall Results
Table 1 and Figure 2 present the performance comparison between Lens and recent multilingual enhancement baselines on multilingual understanding and generation benchmarks, under bilingual and multilingual configurations, respectively. For additional results, including those on Phi-3.5-mini-instruct and multilingual configuration for the other two backbones, please see Appendix E. From the outcomes across all backbones, we have drawn the following key insights:
Lens succeed to achieve a comprehensive improvement for the multilingual capabilities of (M)LLMs without sacrificing original central language performance. Specifically, it enhances the multilingual capabilities of the backbone on both multilingual understanding and generation benchmarks, showing a marked increase in language fidelity during multilingual generation. This effectively mitigates the off-target issue. Moreover, Lens is the only approach that enhances multilingual performance across all languages. In contrast, baseline methods primarily focus on boosting multilingual understanding with little to no improvement in generation tasks. Additionally, methods like QAlign and SDRRL, which rely on translation-based training for multilingual alignment, fall short in effectively enhancing large models’ overall multilingual performance. This suggests that aligning multilingual representations alone is insufficient for fully optimizing multilingual capabilities (Hua et al., 2024). Finally, Lens safeguards the central language from catastrophic forgetting, allowing the resulting model to effectively serve users from diverse linguistic backgrounds.
@@
Using the central language representations within the backbone as a supervision signal proves more effective than relying on external data for supervision. The key distinction between Lens and baseline methods lies in how multilingual performance is enhanced: Lens relies on the model’s internal representation of the central language, while baseline methods depend on external data. This difference make baselines not only fail to improve the target languages but also lead to performance degradation. This phenomenon becomes more pronounced in xSFT-Full when trained with more data. However, the Aya Dataset and Bactrain-X datasets we used are already considered high-quality multilingual resources, widely employed and proven effective in boosting multilingual capabilities in previous models such as mT5 and LLaMA-2 (Li et al., 2023a; Üstün et al., 2024). This highlights that for current extensively trained LLMs such as LLaMA-3 (which has been trained on over 15T data), an over-reliance on external supervision signals may fall short of scalability needs (Cao et al., 2024). We hope Lens could inspire further research to explore more efficient, scalable, and automated supervision signals for multilingual enhancement of the most advanced LLMs.
Figure 2: Results on the multilingual generation benchmark with LLaMA-3-8B-Instruct backbone under the multilingual setting. GPT-4o ratings (on a scale of 1 to 10) are provided for MT-Bench.
@@
Figure 3: The ablation results to verify the effectiveness and impact of different optimization objectives in LSM. MU Performance stands for the average performance on all multilingual understanding benchmarks, while MU Performance is the results on MT-Bench. LF represents language fidelity.
@@ 5.1 Ablation Study
We further conduct ablation studies to demonstrate the effectiveness of our proposed three optimization objectives in LSM. The overall results under the bilingual enhancement setting with LLaMA-3-8B-Instruct backbone are shown in Figure 3. We can draw the following key findings.
@@
The alignment of multiple languages within language-agnostic subspaces mainly impacts multilingual comprehension rather than generation capabilities. As we incrementally raise the coefficient $\lambda_{1}$ of the multilingual alignment objective $\mathcal{L}_{1}$ from zero, Chinese comprehension improves, but its generation ability stays largely unaffected.
Enhancing the separation between representations of different languages in the language-specific subspace is vital for boosting multilingual performance. In particular, as illustrated in the middle part of Figure 3, both comprehension and generation abilities in Chinese improve significantly as the coefficient $\lambda_{l}$ increases. This finding indicates that the commonly accepted notion of merely aligning languages to enhance multilingual capabilities (Cao et al., 2020; Zhu et al., 2023; Li et al., 2024; Hua et al., 2024) may not be sufficient for fully optimizing the multilingual performance of current LLMs. We hope this result encourages future research to focus more on eliciting and leveraging language-specific information within LLMs.
@@
Maintaining the representations of the central language without significant changes can provide a stable and reliable alignment supervision for the target language to be enhanced. As illustrated on the right side of Figure 3, removing the objective to retain English representations leads to a significant decline in the backbone’s Chinese performance. This could be due to alterations in the English representations during optimization, which may cause misalignment in the target for Chinese, thus impacting its performance. However, regarding the English capability, since our modifications are applied to the upper layers of the backbone (layers 31 and 32 in LLaMA-3-8B-Instruct), most of the parameters remain frozen and unaffected. As a result, even without the retaining objective, the backbone’s English ability does not suffer from significant catastrophic forgetting. In the analysis shown in Figure 5, we observe that as the number of updated layers increases, the backbone’s English understanding and generation capabilities are also not impacted by catastrophic forgetting. This can manifest the effectiveness of the retain optimization objective in safeguarding the performance of the central language.
5.2 Impact of Varying the Number of Manipulated Layers
Recent studies on the interpretability of LLMs has sought to reveal the mechanisms underlying their multilingual capabilities (Zhao et al., 2024b; Zhong et al., 2024). A growing consensus suggests that language-specific parameters or neurons are primarily concentrated in the upper layers of these models, while the middle layers tend to process inputs from various languages using a shared and language-agnostic mechanism (Chen et al., 2023a; Wendler et al., 2024; Tang et al., 2024; Kojima et al., 2024; Zhang et al., 2024d). Drawing inspiration from this, our main experiments focus on performing updates solely within the upper layers of the backbone, resulting in a notable improvement in multilingual performance. In Figure 5, we explore the effect of increasing the number of layers involved on the model’s multilingual enhancement. The horizontal axis represents the starting point of the layers where manipulation is applied, with the default endpoint being the final layer. This experiment is performed under the bilingual enhancement with LLaMA-3-8B-Instruct.
It is important to note that modifying only the final layer does not significantly improve either multilingual understanding or generation. This is because language-specific information is not sufficiently enhanced, causing the model to suffer from off-target issues and struggle to represent specific languages accurately. The lack of improvement in multilingual understanding aligns with the findings in Section 5.1, which highlight the critical role of supervision provided by the Push loss ($\mathcal{L}_{2}$).
Our proposed Lens further validates the conclusions of existing works on LLM interpretability and applies these findings to multilingual enhancement of LLMs.
Figure 4: The impact of manipulating different backbone layers on multilingual performance enhancement.
Figure 5: Comparison between our bilingual enhanced model with Chinese-LLaMA series.
5.3 Comparison with Open-Sourced Multilingual-Enhanced LLMs
In Section 4.4, our main experiment primarily compares with reproducible baseline methods for multilingual enhancement. Additionally, we extend our comparisons to several open-source LLMs from the community that leverage private datasets and large-scale post-training to improve multilingual performance. In particular, we focus on the Chinese-LLaMA-3 series, which builds on LLaMA-3 series to enhance Chinese capabilities and includes three different versions:
The experimental results and the resource consumption of different methods are presented in Figure 5 and Table 5 in Appendix E, respectively. The resulting model applied with Lens is identical to the one utilized for bilingual enhancement in Table 1. Remarkably, Lens demonstrates more comprehensive enhancement of the Chinese capabilities with extremely low resource overhead compared to these three models. This reinforces our claim that Lens is an efficient and effective approach for boosting the multilingual capabilities of LLMs. Additionally, all the data leveraged by Lens is publicly accessible, which eliminates the need for laboriously gathering extensive high-quality multilingual datasets and makes it easily shareable with the community.
@@ 5.4 Visualization Analysis
To further confirm whether Lens manipulates language representations within different language subspaces as anticipated, we perform a visualization analysis. Specifically, as shown in Figure 6, we perform Principal Component Analysis (PCA) to visualize the projection of multilingual representations in our obtained language-agnostic subspace and the language-specific subspace. Parallel inputs in seven languages are sourced from the MultiQ datasets (Holtermann et al., 2024). The visualization results indicate that representations of different languages converge within a narrow range in the language-agnostic subspace, while forming distinct clusters in the language-specific subspace, supporting our claim. This also highlights the advantages of Lens in delivering transparent, controllable, and interpretable solutions for the multilingual enhancements of LLMs.

### v2

@@ 3.1 Experimental Setup
We select three English-centric LLMs: LLaMA-3-8B-Instruct (AI@Meta, 2024), LLaMA-3.1-8B-Instruct (Dubey et al., 2024) and Phi-3.5-mini-instruct (Abdin et al., 2024), to fully validate the efficacy and scaleability of our Lens. Notably, although the latter two models undergo more extensive multilingual alignment training, English continues to exhibit the strongest performance.
@@
In the bilingual setting, English (En) serves as the central language, while Chinese (Zh) is the target language for expansion. Chinese is selected due to its growing prominence in the academic focus on multilingual enhancement for LLMs.
In the multilingual setting, we select six target languages, reflecting diverse linguistic families and resource levels. The high-resource languages are Chinese (Zh) and Japanese (Jp); the medium-resource languages are Korean (Ko) and Arabic (Ar); and the low-resource languages are Bengali (Bn) and Swahili (Sw), with English (En) continuing to serve as the central language.
It is important to note that these target languages are classified as out-of-scope in the official model card of the above LLMs, which further underscores their relevance for enhancement.
@@
We sample 300 data points from Aya Dataset (Üstün et al., 2024) for each language to probe the language space and 200 data points from the Bactrian-X dataset (Li et al., 2023a) per language to manipulate the language space. Please see Appendix B for more details of these datasets.
@@
Multilingual Understanding: (1) XCOPA (Ponti et al., 2020), (2) XWinograd (Muennighoff et al., 2023), (3) XStoryCloze (Lin et al., 2022) and (4) M-MMLU (Hendrycks et al., 2021; Lai et al., 2023). Accuracy is the evaluation metric and we randomly sample up to 1,000 data points from each benchmark for evaluation.
Multilingual Generation: (5) MT-Bench (Zheng et al., 2023): A benchmark for open-ended generation to evaluate a model’s ability to follow multi-turn instructions. The evaluation follows the LLM-as-a-judge approach, where GPT-4o is prompted to assign a score directly to a single response on a scale of 1 to 10. It is essential to highlight that the languages targeted for enhancement, as mentioned above, are all within the capability range of GPT-4o, especially given that its official model card (OpenAI, 2024) emphasizes support for low-resource languages such as Swahili (Sw) and Bengali (Bn). This underscores the validity and reliability of the evaluation approach. We employ Language Fidelity (Holtermann et al., 2024) as a metric to assess the consistency between input and output languages, offering a clear measure of how effectively off-target issues are mitigated.
@@ 3.2 Baseline Methods
For comparison, we consider the following baseline methods that enhance LLMs’ multilingual capabilities using multilingual instruction fine-tuning technique: (1) xSFT & xSFT-Full (Ouyang et al., 2022), (2) QAlign (Zhu et al., 2024a), (3) SDRRL (Zhang et al., 2024c) and (4) CLA (Li et al., 2024).
Please refer to Appendix D.1 for the detailed description of the baseline methods.
@@
Table 1: Detailed results on the multilingual understanding and multilingual generation benchmarks under the bilingual setting (English and Chinese). The values in parentheses represent language fidelity. Results highlighted in green indicate an improvement or performance of central language comparable (within a gap of 0.5) to the original backbone, while those highlighted in red signal a decline in performance relative to the original backbone. The best and second-best results in our method and baselines are in bold and underlined, respectively.
@@ 3.3 Implementation Details
Our experiments are implemented with PyTorch (Paszke et al., 2019) and Transformer library (Wolf et al., 2020) on a single NVIDIA A800-SXM4-80GB GPU. The training duration is set to one epoch with the learning rate of 1e-5 and batch size of 8 across all backbones. For more detailed settings, please refer to the Appendix E.
@@ 3.4 Overall Results
Table 1 present the performance comparison between Lens and recent multilingual enhancement baseline methods under bilingual settings. For additional results under multilingual configuration across all three backbones, please see Appendix F.1. From the results across all backbones, we have drawn the following key insights:
Lens achieves comprehensive improvements across diverse languages with different resource levels. It outperforms baselines on both multilingual understanding and generation benchmarks, effectively mitigating the off-target issue. In contrast, baselines primarily favor multilingual understanding, offering minimal or no improvement in generation tasks. In some cases, they compromise language fidelity, exacerbating the off-target problem. Finally, Lens safeguards the central language from catastrophic forgetting. We also compare Lens with LoRA-based SFT, as recent work suggests that it can effectively prevent catastrophic forgetting (Huang et al., 2024). Detailed results and discussions are provided in Appendix F.2.
@@
Using the central language representations within the backbone as a supervision signal proves more effective, efficient and scalable. The key distinction between Lens and baseline methods lies in how multilingual performance is enhanced: Lens relies on the model’s internal representation of the central language, while baseline methods depend on external data. This difference make baselines not only fail to improve the target languages but also lead to performance degradation. However, the Aya Dataset and Bactrain-X datasets we used are already considered high-quality multilingual resources, widely employed and proven effective in boosting multilingual capabilities in previous models such as mT5 and LLaMA-2 (Li et al., 2023a; Üstün et al., 2024). This highlights that for current extensively trained LLMs such as LLaMA-3 (which has been trained on over 15T data), an over-reliance on external supervision signals may fall short of scalability needs (Cao et al., 2024). In addition, the results in Table 6 of Appendix F demonstrate that Lens achieves the best performance with minimal computational overhead.
@@
Figure 2: The ablation results to verify the effectiveness and impact of different optimization objectives in LSM. MU Performance stands for the average performance on all multilingual understanding benchmarks, while MG Performance is the results on MT-Bench. LF represents language fidelity.
Zh Jp Ar Ko Bn Sw
COMET scores on X-to-English Tasks
LLaMA-3 85.40 86.15 84.77 86.07 85.51 78.15
xSFT 70.41 72.40 67.09 72.43 59.52 73.56
QAlign 85.52 85.26 83.11 84.96 83.13 73.66
SDRRL 44.78 45.73 40.87 45.29 45.05 41.51
CLA 85.28 85.35 85.11 85.15 84.36 77.84
Lens 85.64 86.23 85.15 86.07 85.67 80.05
COMET scores on English-to-X Tasks
LLaMA-3 85.28 88.32 76.51 84.53 80.14 71.44
xSFT 83.78 82.22 74.30 81.08 73.40 58.48
QAlign 61.65 58.66 49.41 57.16 41.10 50.96
SDRRL 62.52 57.65 43.83 64.11 68.74 60.00
CLA 80.19 85.75 72.50 82.45 56.73 55.36
Lens 85.59 88.47 79.52 85.77 80.20 71.88
Table 2: Results of baseline methods and our Lens on FLORES-101 benchmark.
@@ 4.1 Ablation Study
We conduct ablation studies to validate the effectiveness of the three optimization objectives in LSM. The results under the bilingual enhancement setting with LLaMA-3-8B-Instruct are shown in Figure 2, leading to the following key findings:
@@
Aligning multiple languages in language-agnostic subspaces primarily enhances multilingual comprehension rather than generation. As the coefficient $\lambda_{1}$ of the alignment loss $\mathcal{L}_{1}$ increases, Chinese comprehension improves, while its generation ability remains slightly unaffected.
Separating language representations in language-specific subspaces is crucial for multilingual performance. As shown in the middle of Figure 2, increasing $\lambda_{l}$ significantly boosts both comprehension and generation in Chinese, yielding greater benefits than mere language alignment. This suggests that the conventional approach of aligning languages (Cao et al., 2020; Zhu et al., 2023; Li et al., 2024; Hua et al., 2024) may be insufficient for fully optimizing multilingual LLMs. Further discussion in Appendix F.3 elaborates on this insight.
To further illustrate the benefit of the distinction between representations of different languages within the language-specific subspace, we conduct additional experiments on the multilingual machine translation task and results are shown in Table 2. The consistent performance improvements achieved by Lens confirm that it captures the subtle linguistic nuances required for precise and fluent translations. Detailed experimental setups and analysis are provided in Appendix F.4.
@@
Maintaining stable representations of the central language provides reliable alignment supervision. As shown in the rightmost part of Figure 2, removing the retention objective for English leads to a significant drop in Chinese performance, likely due to misalignment caused by changes in English representations. However, since our modifications only affect the upper layers (layers 31 and 32) of LLaMA-3-8B-Instruct, most parameters remain frozen, preventing catastrophic forgetting of English capabilities. Further analysis in §4.2 confirms that increasing the number of updated layers does not negatively impact English performance, highlighting the efficacy of the retention objective.
@@
4.2 Impact of Varying the Number of Manipulated Layers and Training Data
Building on recent studies on LLM interpretability (Zhao et al., 2024c; Zhong et al., 2024), we focus on updating only the upper layers of the backbone. Figure LABEL:subfig:layer_analysis examines the impact of modifying different layers under the bilingual enhancement setting with LLaMA-3-8B-Instruct. The horizontal axis denotes the starting layer for updates, with the final layer as the default endpoint.
Intermediate-layer processing in English benefits multilingual understanding. If the target language is partitioned into the language-specific subspace too early, it weakens multilingual comprehension. However, modifying only the final layer is insufficient for enhancing either understanding or generation, as language-specific information remains underdeveloped, leading to off-target errors.
Lens further validates the conclusions of existing works on LLM interpretability and applies these findings to multilingual enhancement. Please refer to Appendix F.5 for more analysis.
Results in Figure LABEL:subfig:data_num indicate that increasing the amount of training data leads to diminishing returns for Lens. This finding reinforces our claim that for extensively pre-trained LLMs such as LLaMA-3 (trained on over 15T tokens), over-reliance on more training data falls short of meeting scalability needs. Instead of focusing on larger training datasets, it is more critical to identify supervision signals that are both reliable and scalable. This directly motivates us to seek internal supervision from the central language with the backbone itself.
4.3 Comparison with Open-Source Multilingual-Enhanced LLMs
In §3.4, we compare with reproducible baselines. Additionally, we extend our comparisons to open-source LLMs that leverage private datasets and large-scale post-training to improve multilingual performance. In particular, we focus on the Chinese-LLaMA-3 series, which builds on LLaMA-3 series to enhance Chinese capabilities and includes three versions. For a detailed description of these models, please refer to Appendix D.2.
Results and the resource consumption of different methods are presented in Figure LABEL:subfig:comprare_chinese and Table 6 in Appendix F, respectively. The resulting model applied with Lens is identical to the one utilized for bilingual enhancement in Table 1. Lens demonstrates more comprehensive enhancement of the Chinese capabilities with extremely low resource overhead compared to these three models. Also, all the data leveraged by Lens is publicly accessible, which eliminates the need for laboriously gathering extensive high-quality multilingual datasets and makes it easily shareable with the community.
@@ 4.4 Visualization Analysis
To further confirm whether Lens manipulates language representations within different language subspaces as anticipated, as shown in Figure 5, we perform Principal Component Analysis (PCA) to visualize the projection of multilingual representations in the language-agnostic and -specific subspace. Parallel inputs in seven languages are sourced from the MultiQ datasets (Holtermann et al., 2024). Results indicate that representations of different languages converge within a narrow range in the language-agnostic subspace, while forming distinct clusters in the language-specific subspace, supporting our claim. This also highlights the advantages of Lens in delivering transparent, controllable, and interpretable solutions for the multilingual enhancements of LLMs.
4.5 Impact of Language Proximity
In §3.4, we focus on languages that are out-of-scope and under-represented in the backbone, selected based on their typological and script differences from the central language, English. To further highlight the generalizability of Lens, we expand our analysis to include three additional languages: Spanish, French, and German. These languages, being typologically closer to English, are already better represented in the backbone model. As illustrated in Table 7 of Appendix F.6, Lens still exhibit notable improvements. For a more comprehensive discussion, please refer to Appendix F.6.
5 Related Works
Multilingual Large Language Model
Pretraining on multilingual data is a common approach to gain multilingual capabilities (Conneau and Lample, 2019; Xue et al., 2020; Lin et al., 2022; Shliazhko et al., 2022; Wei et al., 2023; Xue et al., 2022; Le Scao et al., 2023; Blevins et al., 2024). However, due to the uneven distribution of data in pretraining corpora, current LLMs or MLLMs exhibit uneven language capabilities, with most models heavily biased towards English (Jiang et al., 2023; AI@Meta, 2024; Abdin et al., 2024). Moreover, pretraining from scratch is computationally intensive. These limitations have directly sparked research into expanding or enhancing the language capabilities of current LLMs or MLLMs.
Multilingual Enhancement for LLMs
The former focuses on leveraging the LLMs’ own translation capabilities to translate low-resource language inputs into the central language, and then generating a response (Shi et al., 2023; Huang et al., 2023; Qin et al., 2023; Etxaniz et al., 2024; Zhang et al., 2024b). However, Liu et al. (2024a) reveal the limitations of these methods, showing they are not optimal for real-world scenarios and highlighting the necessity of more comprehensive multilingual enhancement.
The latter aims to conduct further multilingual post-training to inject language knowledge or achieve cross-lingual alignment, including ways of continual pre-training (Zhang et al., 2021b; Cui et al., 2023; Chen et al., 2023b; Lin et al., 2024b; Kuulmets et al., 2024; Jaavid et al., 2024) and instruction tuning (Muennighoff et al., 2023; Chen et al., 2023c; Indurthi et al., 2024; Ahuja et al., 2024; Lai and Nissim, 2024; Zhang et al., 2024c; Zhu et al., 2024a; Li et al., 2024; Zhao et al., 2024d). For example, Cui et al. (2023) attempt to inject Chinese knowledge into LLaMA by conducting continual pre-training on a large-scale Chinese corpus, while Zhu et al. (2023) focus more on building language alignment through cross-lingual instruction tuning and translation training.
Our proposed Lens stands out from existing methods in that we seek multilingual supervision signals from the internal language representation space of the LLMs, rather than relying heavily on external multilingual datasets as in the above methods, which offers fresh insights and new opportunities for enhancing the multilingual capabilities of LLMs both efficiently and effectively. For more discussions on the theoretical foundation of language-agnostic and language-specific subspaces within LLMs, please refer to Appendix G.
Representation Engineering
Editing or manipulating representation within LLMs has garnered increasing attention due to its transparency and lightweight properties (Zou et al., 2023). This is theoritically rooted from Linear Representation Hypothesis (Mikolov et al., 2013; Nanda et al., 2023; Park et al., 2024). Building upon this, existing works attempt to edit representations at inference time to develop models that are more truthful (Li et al., 2023b; Campbell et al., 2023; Zhang et al., 2024a), and harmless (Lee et al., 2024; Uppaal et al., 2024). We expand and implement this paradigm for the multilingual enhancement of LLMs by focusing on representations during the training phase, ensuring that the efficiency of LLMs remains unaffected during the inference phase.